Given that the variance is Q, determine the standard deviation in terms of Q.



a) 0.5Q


b) 2Q


c) Q


d) Q2

1
Expert's answer
2022-03-24T14:29:49-0400

a) "\sigma^2=0.5Q \to\sigma =\sqrt{0.5Q}\approx 0.71\sqrt{Q}."

b) "\sigma^2=2Q \to\sigma =\sqrt{2Q}\approx 1.41\sqrt{Q}."

c) "\sigma^2=Q \to\sigma =\sqrt{Q}."

d) "\sigma^2=Q^2 \to\sigma =\sqrt{Q^2}=Q."